During 2020, Sarasota Furniture Company purchases a carload of wicker chairs. The manufacturer sells the chairs to Sarasota for a lump sum of $131,670 because it is discontinuing manufacturing operations and wishes to dispose of its entire stock. Three types of chairs are included in the carload. The three types and the estimated selling price for each are listed below.

Type

No. of Chairs

Estimated Selling
Price Each

Lounge chairs

880 $90

Armchairs

660 80

Straight chairs

1,540 50


During 2020, Sarasota sells 440 lounge chairs, 220 armchairs, and 264 straight chairs.

What is the amount of gross profit realized during 2020? What is the amount of inventory of unsold straight chairs on December 31, 2020? (Round cost per chair to 2 decimal places, e.g. 78.25 and final answer to 0 decimal places, e.g. 5,845.)

Gross profit realized during 2020

$enter a dollar amount

Amount of inventory of unsold straight chairs

$enter a dollar amount
Answer #1

Type

No. of chairs (i)

Price per chair (ii)

Sales value (i) x (ii)

Sales value percentage

Lounge chairs

880

$90

79,200

79,200/209,000 = 37.89473684%

Armchairs

660

80

52,800

52,800/209,000 = 25.26315789%

Straight chairs

1,540

50

77,000

77,000/209,000 = 36.84210526%

Total

209,000

Type

Total cost   

Cost per Chair

Lounge chairs

131,670 x 37.89473684% = 49,896

49,896/880 = 56.7

Armchairs

131,670 x 25.26315789 = 33,264

33,264/660 = 50.4

Straight chairs

131,670 x 36.84210526% = 48,510

48,510/1,540 = 31.5

Gross profit = Sales - Cost of goods sold

= (440 x 90 + 220 x 80 + 264 x 50) - (440 x 56.7 + 220 x 50.4 + 264 x 31.5)

= (39,600 + 17,600 + 13,200) - (24,948 + 11,088 + 8,316)

= 70,400 - 44,352

= $26,048

Ending inventory schedule

Type Number of chairs (i) Cost per chair (ii) Total cost (i) x (ii)
Lounge chairs 440 56.7 24,948
Armchairs 440 50.4 22,176
Straight chairs 1,276 31.5 40,194
Total 87,318

Amount of inventory of unsold straight chairs on December 31, 2020 = $87,318

Kindly comment if you need further assistance. Thanks